# All or None: Identifiable Linear Properties of Next-token Predictors in Language Modeling

### v1

@@ B.5 Proof of Theorem 5
Notice that $\mathrm{rank}({\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{G}}})=\mathrm{dim}(\mathcal{N})$ and $\mathrm{rank}({\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{{\bm{\mathrm{G}}}})=\mathrm{dim}(\tilde{\mathcal{N}})$ by Lemma 17, and it also holds $\mathrm{dim}(\mathcal{N})=\mathrm{dim}(\mathcal{M})=\mathrm{dim}(\tilde{\mathcal{M}})=\mathrm{dim}(\tilde{\mathcal{N}})$ by Lemma 2 (i) and Step 3. Let $k:=\mathrm{dim}(\mathcal{N})$. Using this in 121, we obtain:
@@
This shows that, combined with 120 we have $k\leq\mathrm{rank}({\bm{\mathrm{N}}})\leq k$, which means that $\mathrm{rank}({\bm{\mathrm{N}}})=k$. Taking Steps 2, 3, 4, 5 together, we have that:
@@
where in the last line we included the expression for the conditional probability $p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y\mid{\bm{\mathrm{x}}})=\exp({\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{g}}}(y))/Z({\bm{\mathrm{x}}})$ from 1, and we denoted $\tilde{Z}({\bm{\mathrm{x}}}):={Z}({\bm{\mathrm{x}}})\big{/}{\exp\big{(}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{g}}}(y_{0})-\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}\tilde{\bm{\mathrm{g}}}(y_{0})\big{)}}$. To obtain the value of $\tilde{Z}$ we consider the sum over all $y\in\mathcal{A}$ for 128, giving:
@@
which means that, from 128 we have:
@@ B.6 Proof of Corollary 6
To highlight deviations, we present a proof that follows a somewhat analogous argument to the proof of Roeder et al. [3]; a direct proof may show ${\bm{\mathrm{N}}}={\bm{\mathrm{M}}}^{-1}$ relying on Theorem 5. We relax condition 2 and use the fact the assumption that $\mathrm{SIm}({\bm{\mathrm{f}}})=\mathbb{R}^{d}$, which is a milder condition to requiring that $\mathrm{SIm}({\bm{\mathrm{f}}}_{0})=\mathbb{R}^{d}$. We prove the following:
@@
Our proof follows a similar technique to Lachapelle et al. [25, Theorem B.4].
@@
Identifiability of ${\bm{\mathrm{g}}}_{0}$. Next, we consider the implication for ${\bm{\mathrm{g}}}_{0}$. We start again from the pivot difference of Equation 140:
@@
where we substituted $\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})={\bm{\mathrm{M}}}^{-1}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})$ from 95. Therefore, taking $d$ points ${\bm{\mathrm{x}}}\in\mathsf{Seq}(\mathcal{A})$, such that the matrix ${\bm{\mathrm{F}}}$ is invertible, restarting from the transpose of 149 we obtain:
@@ C.1 Relational Linear Steering Property
We here want to discuss an additional linear property besides those presented in Section 4, termed linear steering property. This behavior is also referred to as the linear intervening property by Park et al. [23].
It has been empirically observed that there exist steering vectors that influence next-token predictions [23, 2, 61, 62], in the following sense: If ${\bm{\mathrm{v}}}$ encode the average difference between English to Italian embeddings, adding ${\bm{\mathrm{v}}}$ to ${\bm{\mathrm{f}}}({\bm{\mathrm{s}}})$ for the sentence ${\bm{\mathrm{s}}}=$The king sits on the” would change the most-likely next token prediction $y=$throne” to $\hat{y}=$trono”, and similarly this applies for other sentences, affecting the most-likely next-token prediction to move from the English token to the Italian counterpart.
@@ C.2 Proof of Lemma 9
and multiplying both sides of (164) from the left with the pseudo-inverse of ${\bm{\mathrm{F}}}^{\top}$ we get:
@@
This shows the claim.
@@
In this context, we exclude the case $\beta=0$ because, for the zero vector ${\bm{\mathrm{a}}}=0$, a contradiction arises. Specifically, any vector ${\bm{\mathrm{b}}}\in\mathbb{R}^{d}$ would be trivially parallel to ${\bm{\mathrm{a}}}$ (i.e., ${\bm{\mathrm{a}}}=\beta{\bm{\mathrm{b}}}$ with $\beta=0$). However, conversely, we would obtain that no scalar $\beta\in\mathbb{R}$ exists such that ${\bm{\mathrm{b}}}=\beta{\bm{\mathrm{a}}}$.
@@ C.4 Proof of Proposition 13
we can then take any ${\bm{\mathrm{g}}}_{j}(y_{i})\in\Gamma$, and multiply their transpose times both sides of (185) from the right. We then get
@@
We can then substitute the expression on the RHS in (188) to (184) to obtain:
@@
$=\exp\big{(}{{\bm{\mathrm{g}}}_{j}(y_{i})^{\top}{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+{\bm{\mathrm{g}}}_{j}(y_{i})^{\top}{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}}}\big{)}\,\frac{e^{({\bm{\mathrm{g}}}(y_{j})^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})}}{Z({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}};\mathcal{Y}_{P})}$ (Use 188)
@@
Denote with ${\bm{\mathrm{w}}}{{}_{i}}:={\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}^{\top}{\bm{\mathrm{g}}}(y_{i})$ and with $b_{i}:={\bm{\mathrm{g}}}(y_{i})^{\top}{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}}$. Then using this and 193 inside 191 we get
@@
$=\frac{\exp\big{(}{\bm{\mathrm{w}}}_{i}^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+b_{i}\big{)}}{\sum_{y_{i}\in\mathcal{Y}_{P}}\exp\big{(}{\bm{\mathrm{w}}}_{i}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+b_{i}\big{)}}$ (Substitute for ${\bm{\mathrm{w}}}_{i}$ and $b_{i}$, and 193)
@@ D.1 Proof of Proposition 14
where in the last line we used that ${\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{P}}}_{\mathcal{F}}={\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}}={\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{P}}}_{\mathcal{M}}={\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{P}}}_{\mathcal{M}}$ by Lemma 2 (vi). We now substitute the expression for $(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})$ based on the RHS of the equivalence relation(9) in 197 to get
@@
where in the last line we multiplied on the left by the orthogonal projector ${\bm{\mathrm{P}}}_{\Gamma}$ and used that ${\bm{\mathrm{M}}}={\bm{\mathrm{M}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}$. We can now show that we can substitute 200 into the left-hand side of 198:
@@
Step 2. We proceed to show that $\tilde{\Gamma}\subseteq\tilde{\mathcal{N}}$. First, we have to show that ${\bm{\mathrm{P}}}_{\tilde{\Gamma}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}={\bm{\mathrm{P}}}_{\tilde{\Gamma}}$. To this end, notice that $({\bm{\mathrm{N}}}^{\top})^{+}:\tilde{\mathcal{N}}\to\mathcal{N}$. Therefore, we have from Equation 203:
@@
$=\mathrm{Im}({\bm{\mathrm{N}}}^{+}{\bm{\mathrm{P}}}_{\Gamma}^{\top})$ ($\mathrm{ker}({\bm{\mathrm{A}}})^{\perp}=\mathrm{Im}({\bm{\mathrm{A}}}^{\top})$, for any matrix ${\bm{\mathrm{A}}}$ [60])
@@
What happens if $\Gamma_{\bm{\mathrm{q}}}\not\subseteq\mathcal{M}$?. We discuss the case when the condition in Proposition 14 is not met due to $\Gamma_{\bm{\mathrm{q}}}:=\mathrm{Im}({\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}^{\top}{\bm{\mathrm{P}}}_{\Gamma})\not\subseteq\mathcal{M}$. We show that even if a model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ linearly represents ${\bm{\mathrm{q}}}$ on $\Gamma$, a $\sim_{EL}$-equivalent model $(\tilde{{\bm{\mathrm{f}}}},\tilde{\bm{\mathrm{g}}})$ may not. This is due to the fact that the information contained in $\mathcal{F}\setminus\mathcal{M}$, used to relationally represent ${\bm{\mathrm{q}}}$, may not be linearly transformed on another $\sim_{EL}$-equivalent model.
@@
We take the first term on the RHS of (217): by inserting the projector ${\bm{\mathrm{P}}}_{\Gamma_{\bm{\mathrm{q}}}}$, we rewrite it as follows:
@@
where we used the identity ${\bm{\mathrm{I}}}={\bm{\mathrm{P}}}_{\mathcal{M}}+({\bm{\mathrm{I}}}-{\bm{\mathrm{P}}}_{\mathcal{M}})$ to separate the contributions, inside and outside $\mathcal{M}$. We can thus rewrite 217 as:
@@
where the first term follows from the equivalence relation 9, i.e., ${\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})={\bm{\mathrm{M}}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})$, and we used $\tilde{\bm{\mathrm{f}}}^{2}({\bm{\mathrm{x}}})=(\tilde{f}_{1}({\bm{\mathrm{x}}})^{2},\ldots,\tilde{f}_{\tilde{d}}({\bm{\mathrm{x}}})^{2})$ to denote the square of the components of $\tilde{{\bm{\mathrm{f}}}}$. Notice that, this choice is allowed since the components of ${\bm{\mathrm{f}}}$ outside $\mathcal{M}$, i.e., those in $\mathcal{F}\setminus\mathcal{M}$, can be arbitrarily chosen and do not contribute to the dot-product with ${\bm{\mathrm{g}}}_{0}$. Therefore, substituting this expression to 219 we get:
@@
and substituting ${\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})={\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{N}}}^{+}\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})$, implied by the LHS of 197 and from the equality in 200, we have that:

### v2

@@ B.6 Proof of Theorem 5
Notice that $\mathrm{rank}({\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{G}}})=\mathrm{dim}(\mathcal{N})$ and $\mathrm{rank}({\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{{\bm{\mathrm{G}}}})=\mathrm{dim}(\tilde{\mathcal{N}})$ by Lemma 17, and it also holds $\mathrm{dim}(\mathcal{N})=\mathrm{dim}(\mathcal{M})=\mathrm{dim}(\tilde{\mathcal{M}})=\mathrm{dim}(\tilde{\mathcal{N}})$ by Lemma 2 (i) and Step 3. Let $k:=\mathrm{dim}(\mathcal{N})$. Using this in Equation (125), we obtain:
@@
This shows that, combined with Equation (124) we have $k\leq\mathrm{rank}({\bm{\mathrm{N}}})\leq k$, which means that $\mathrm{rank}({\bm{\mathrm{N}}})=k$. Taking Steps 2, 3, 4, 5 together, we have that:
@@
where in the last line we included the expression for the conditional probability $p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y\mid{\bm{\mathrm{x}}})=\exp({\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{g}}}(y))/Z({\bm{\mathrm{x}}})$ from Equation (1), and we denoted $\tilde{Z}({\bm{\mathrm{x}}}):={Z}({\bm{\mathrm{x}}})\big{/}{\exp\big{(}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{g}}}(y_{0})-\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}\tilde{\bm{\mathrm{g}}}(y_{0})\big{)}}$. To obtain the value of $\tilde{Z}$ we consider the sum over all $y\in\mathcal{A}$ for Equation (132), giving:
@@
which means that, from Equation (132) we have:
@@ B.7 Proof of Corollary 6
To highlight deviations, we present a proof that follows a somewhat analogous argument to the proof of [Roeder et al., 2021]; a direct proof may show ${\bm{\mathrm{N}}}={\bm{\mathrm{M}}}^{-1}$ relying on Theorem 5. We relax condition 2 and use the fact the assumption that $\mathrm{SIm}({\bm{\mathrm{f}}})=\mathbb{R}^{d}$, which is a milder condition to requiring that $\mathrm{SIm}({\bm{\mathrm{f}}}_{0})=\mathbb{R}^{d}$. We prove the following:
@@
Our proof follows a similar technique to Lachapelle et al. [2023, Theorem B.4].
@@
Identifiability of ${\bm{\mathrm{g}}}_{0}$. Next, we consider the implication for ${\bm{\mathrm{g}}}_{0}$. We start again from the pivot difference of Equation Equation (144):
@@
where we substituted $\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})={\bm{\mathrm{M}}}^{-1}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})$ from Equation (99). Therefore, taking $d$ points ${\bm{\mathrm{x}}}\in\mathsf{Seq}(\mathcal{A})$, such that the matrix ${\bm{\mathrm{F}}}$ is invertible, restarting from the transpose of Equation (153) we obtain:
@@ C.1 Relational Linear Steering Property
We here want to discuss an additional linear property besides those presented in Section 4, termed linear steering property. This behavior is also referred to as the linear intervening property by Park et al. [2024a].
It has been empirically observed that there exist steering vectors that influence next-token predictions [Park et al., 2024a, Hernandez et al., 2024, Hase et al., 2024, Arditi et al., 2024], in the following sense: If ${\bm{\mathrm{v}}}$ encode the average difference between English to Italian embeddings, adding ${\bm{\mathrm{v}}}$ to ${\bm{\mathrm{f}}}({\bm{\mathrm{s}}})$ for the sentence ${\bm{\mathrm{s}}}=$The king sits on the” would change the most-likely next token prediction $y=$throne” to $\hat{y}=$trono”, and similarly this applies for other sentences, affecting the most-likely next-token prediction to move from the English token to the Italian counterpart.
@@ C.2 Proof of Lemma 8
and multiplying both sides of Equation (168) from the left with the pseudo-inverse of ${\bm{\mathrm{F}}}^{\top}$ we get:
@@
This shows the claim. ∎
@@
We exclude the case $\beta=0$ because, for ${\bm{\mathrm{a}}}=0$, a contradiction arises. Specifically, any vector ${\bm{\mathrm{b}}}\in\mathbb{R}^{d}$ would be trivially parallel to ${\bm{\mathrm{a}}}$ (i.e., ${\bm{\mathrm{a}}}=\beta{\bm{\mathrm{b}}}$ with $\beta=0$). However, conversely, we would obtain that no scalar $\beta\in\mathbb{R}$ exists such that ${\bm{\mathrm{b}}}=\beta{\bm{\mathrm{a}}}$.
@@ C.4 Proof of Proposition 13
we can then take any ${\bm{\mathrm{g}}}_{j}(y_{i})\in\Gamma$, and multiply their transpose times both sides of Equation (189) from the right. We then get
@@
We can then substitute the expression on the RHS in Equation (192) to Equation (188) to obtain:
@@
$=\exp\big{(}{{\bm{\mathrm{g}}}_{j}(y_{i})^{\top}{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+{\bm{\mathrm{g}}}_{j}(y_{i})^{\top}{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}}}\big{)}\,\frac{e^{({\bm{\mathrm{g}}}(y_{j})^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})}}{Z({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}};\mathcal{Y}_{P})}$ (Use Equation (192))
@@
Denote with ${\bm{\mathrm{w}}}{{}_{i}}:={\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}^{\top}{\bm{\mathrm{g}}}(y_{i})$ and with $b_{i}:={\bm{\mathrm{g}}}(y_{i})^{\top}{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}}$. Then using this and Equation (197) inside Equation (195) we get
@@
$=\frac{\exp\big{(}{\bm{\mathrm{w}}}_{i}^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+b_{i}\big{)}}{\sum_{y_{i}\in\mathcal{Y}_{P}}\exp\big{(}{\bm{\mathrm{w}}}_{i}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+b_{i}\big{)}}$ (Substitute for ${\bm{\mathrm{w}}}_{i}$ and $b_{i}$, and Equation (197))
@@ D.1 Proof of Theorem 14
where in the last line we used that ${\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{P}}}_{\mathcal{F}}={\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}}={\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{P}}}_{\mathcal{M}}={\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{P}}}_{\mathcal{M}}$ by Lemma 2 (vi). We now substitute the expression for $(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})$ based on the RHS of the equivalence relation Equation (9) in Equation (201) to get
@@
where in the last line we multiplied on the left by the orthogonal projector ${\bm{\mathrm{P}}}_{\Gamma}$ and used that ${\bm{\mathrm{M}}}={\bm{\mathrm{M}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}$. We can now show that we can substitute Equation (204) into the left-hand side of Equation (202):
@@
Step 2. We proceed to show that $\tilde{\Gamma}\subseteq\tilde{\mathcal{N}}$. First, we have to show that ${\bm{\mathrm{P}}}_{\tilde{\Gamma}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}={\bm{\mathrm{P}}}_{\tilde{\Gamma}}$. To this end, notice that $({\bm{\mathrm{N}}}^{\top})^{+}:\tilde{\mathcal{N}}\to\mathcal{N}$. Therefore, we have from Equation (207):
@@
$=\mathrm{Im}({\bm{\mathrm{N}}}^{+}{\bm{\mathrm{P}}}_{\Gamma}^{\top})$ ($\mathrm{ker}({\bm{\mathrm{A}}})^{\perp}=\mathrm{Im}({\bm{\mathrm{A}}}^{\top})$, for any matrix ${\bm{\mathrm{A}}}$ [Axler, 2015])
@@
What happens if $\Gamma_{\bm{\mathrm{q}}}\not\subseteq\mathcal{M}$?. We discuss the case when the condition in Theorem 14 is not met due to $\Gamma_{\bm{\mathrm{q}}}:=\mathrm{Im}({\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}^{\top}{\bm{\mathrm{P}}}_{\Gamma})\not\subseteq\mathcal{M}$. We show that even if a model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ linearly represents ${\bm{\mathrm{q}}}$ on $\Gamma$, a $\sim_{EL}$-equivalent model $(\tilde{{\bm{\mathrm{f}}}},\tilde{\bm{\mathrm{g}}})$ may not. This is due to the fact that the information contained in $\mathcal{F}\setminus\mathcal{M}$, used to relationally represent ${\bm{\mathrm{q}}}$, may not be linearly transformed on another $\sim_{EL}$-equivalent model.
@@
We take the first term on the RHS of Equation (221): by inserting the projector ${\bm{\mathrm{P}}}_{\Gamma_{\bm{\mathrm{q}}}}$, we rewrite it as follows:
@@
where we used the identity ${\bm{\mathrm{I}}}={\bm{\mathrm{P}}}_{\mathcal{M}}+({\bm{\mathrm{I}}}-{\bm{\mathrm{P}}}_{\mathcal{M}})$ to separate the contributions, inside and outside $\mathcal{M}$. We can thus rewrite Equation (221) as:
@@
where the first term follows from the equivalence relation Equation (9), i.e., ${\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})={\bm{\mathrm{M}}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})$, and we used $\tilde{\bm{\mathrm{f}}}^{2}({\bm{\mathrm{x}}})=(\tilde{f}_{1}({\bm{\mathrm{x}}})^{2},\ldots,\tilde{f}_{\tilde{d}}({\bm{\mathrm{x}}})^{2})$ to denote the square of the components of $\tilde{{\bm{\mathrm{f}}}}$. Notice that, this choice is allowed since the components of ${\bm{\mathrm{f}}}$ outside $\mathcal{M}$, i.e., those in $\mathcal{F}\setminus\mathcal{M}$, can be arbitrarily chosen and do not contribute to the dot-product with ${\bm{\mathrm{g}}}_{0}$. Therefore, substituting this expression to Equation (223) we get:
@@
and substituting ${\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})={\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{N}}}^{+}\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})$, implied by the LHS of Equation (201) and from the equality in Equation (204), we have that: